OpenStudy (anonymous):

(a) At what height above the earth is the acceleration due to gravity 44% of its value at the surface? m (b) What is the speed of a satellite orbiting at that height? m/s

OpenStudy (anonymous):

\[F_{grav} = G\frac{mM}{r^2}\] where r is the distance from the CENTER of the earth. Let R be the radius of the earth. \[F_{grav,surface} = G\frac{mM}{R^2}\] \[.44F_{grav,surface} = G\frac{44}{100}\frac{mM}{R^2}=G\frac{mM}{r^2}\] Solve for r

OpenStudy (anonymous):

what is mM

OpenStudy (anonymous):

two masses. One is the mass of the earth, the other is the mass of the object. It doesn't matter which one you use for which, as long as you are consistent.

OpenStudy (shamim):

Dont u need the solution of ur question b)

OpenStudy (anonymous):

this is what i have so far?

OpenStudy (anonymous):

@Vandreigan

OpenStudy (anonymous):

You won't need most of that stuff. \[G(.44)\frac{mM}{R^2} = G\frac{mM}{r^2}\] \[\frac{.44}{R^2}=\frac{1}{r^2}\] \[r = \sqrt{\frac{100R^2}{44}}\] \[r = \frac{5R}{\sqrt{11}} = \frac{5\sqrt{11}R}{11}\]

OpenStudy (anonymous):

wouldn't that just come out to R=r

OpenStudy (anonymous):

would R be 6778100 (the radius of earth in M)

OpenStudy (anonymous):

Nope \[r = \frac{5R}{\sqrt{11}}\] You can just plug in the radius of the earth to find r. Remember, though. This is the distance from the CENTER of the earth.

OpenStudy (anonymous):

i get 10218370.2 (but i don't think thats right)

OpenStudy (anonymous):

i only have like 10 minutes until my computer dies

OpenStudy (anonymous):

R = 6371000m Therefore, r = 9604643.882m

OpenStudy (anonymous):

Distance above the ground is then: r-R = 3233643.882

OpenStudy (anonymous):

ooooohhh

OpenStudy (anonymous):

so how do i find the speed

OpenStudy (anonymous):

As for the speed of the satellite, we can just use: \[a = \frac{v^2}{r}\] Where a is the acceleration due to gravity at the distance r (44% of what it is on the surface)

OpenStudy (anonymous):

6435.46?

OpenStudy (anonymous):

\[v = \sqrt{.44*g*r}\] g = 9.81 m/s^2 r = 9604643.882m

OpenStudy (anonymous):

the answer was 435.46 thanks so much for your help! gotta go